OpenStudy (crashonce):

A,2,B are in AP and A,5,B are in GP. FInd A and B

OpenStudy (anonymous):

yes @CrashOnce

OpenStudy (crashonce):

?

OpenStudy (crashonce):

@hartnn

OpenStudy (crashonce):

@ganeshie8

OpenStudy (crashonce):

@hartnn

hartnn (hartnn):

A,2,B are in AP so, A+B = 2(2) = 4 got this ?

OpenStudy (crashonce):

yes

hartnn (hartnn):

A,5, B are in GP AB = 25 you know these formulas of AP and GP ?

hartnn (hartnn):

x,y,z in AP means 2y = x+z x,y,z in GP means y^2 = xz

OpenStudy (crashonce):

yes

hartnn (hartnn):

you can get a quadratic in A or B

OpenStudy (crashonce):

The second part of the question is: Which of the following quadratics has A and B as its root? 1)x^2+2x+5=0. 2) x^2-5x+2=0. 3)x^2-4x+25=0. 4)x^2+4x+25=0. 5) x^2-25x-4=0

hartnn (hartnn):

AB = 25 A+B = 4

hartnn (hartnn):

you found A,B ?

hartnn (hartnn):

quadratic : \(x^2 - (sum)x + product =0\) \(x^2 - (A+B)x + AB =0\)

OpenStudy (crashonce):

so its 3). ok thanks again

hartnn (hartnn):

yes and welcome ^_^